OpenStudy (anonymous):

Use the iterative rule to find the seventh term of the sequence: an = 3n – 1. 4 minutes ago

OpenStudy (anonymous):

728?

OpenStudy (bloomlocke367):

well do you know what the iterative rule is?

OpenStudy (anonymous):

I tried it again and I got 2187

OpenStudy (bloomlocke367):

I thought you just have to put 7 in for x.... umm.. hang on... let me do some more research

OpenStudy (anonymous):

The Iterative Rule is: an=a1 n-1

OpenStudy (bloomlocke367):

is it supposed to be\[a_n=3^{n-1}\]

OpenStudy (anonymous):

Yes! Correct!

OpenStudy (bloomlocke367):

okay.. so, I still believe you plug 7 in for n.... so you'll have\[a_n=3^{7-1}=3^6\]

OpenStudy (bloomlocke367):

can you do that?

OpenStudy (anonymous):

Could you show me? Im not good at this

OpenStudy (bloomlocke367):

\[3^6=3\times3\times3\times3\times3\times3\]

OpenStudy (bloomlocke367):

and you were close with your first answer.

OpenStudy (anonymous):

3^ 15?

OpenStudy (bloomlocke367):

no..... you said 728... that's really close to the answer..

OpenStudy (anonymous):

2187 so this is my answer

OpenStudy (bloomlocke367):

no, do you have a calculator?

OpenStudy (anonymous):

Sorry bout that my laptop is freezing! yes I do I just got 728 now

OpenStudy (bloomlocke367):

you should be getting 729...

OpenStudy (anonymous):

Sorry. do you have time for a few more

OpenStudy (bloomlocke367):

I'm in a Lesson right now... ummm so not really.. sorry. but no need to apologize! XD

OpenStudy (anonymous):

Ah k

OpenStudy (bloomlocke367):

sorry!